OpenStudy (anonymous):

Using the unit circle what is the exact value of "tan (pi/6)"?

OpenStudy (anonymous):

A sqrt( 3)/2 B. 1/2 C sqrt(3)/3 D sqrt(3)

OpenStudy (anonymous):

tan pi/6 = (sinpi/6)/(cos pi/6)= (1/2) / sqr3/2)=1/sqr3 or sqr3/3

OpenStudy (anonymous):

it's C.